If a firm has retained earnings of $2.6 million, a common shares account of $4.6 million, and additional paid-in capital of $9.2 million, how would these accounts change in response to a 10 percent stock dividend? Assume market value of equity is equal to book value of equity. (Enter your answers in dollars not in millions. Input all amounts as positive values. Indicate the direction of the effect by selecting "increase," "decrease," or "no change" from the drop-down menu.)

Reatained Earnings _____ to _____

Common Stock _______ to _______

Additional paid-in capital ________ to _______
Answer #1

10 % Stock dividend means value of shares in the common stock and additional capital will be increased in form of dividend and the amount increased will be transferred from retaimed earnings.

Dividend is calculated as follows

Common shares = 4600000 * 10 % = $460000

Additional paid in capital = 9200000 * 10% = $920000

New balance

Common shares = 4600000 + 460000 = $5060000

Additional paid in capital = 9200000 + 920000 = $10120000

Retained earning = Balance - ( Dividend on common shares + paid in capital)

= 2600000 - ( 460000 + 920000 )

= $1220000

Retained earning 2.6 milliom to 1.22 million

Common shares 4.6 million to 5.06 million

Additinal capital 9.2 million to 10.12 million